Douglas invests money in two simple interest accounts. He invests three times as much in an account paying 14% as he does in an account paying 5%. If he earns $152.75 in interest in one year from both accounts combined, how much did he invest altogether?
Total Principal in Both Accounts = $

Found 3 solutions by mananth, ikleyn, greenestamps:
Answer by mananth(16946)


Douglas invests money in two simple interest accounts. He invests three times as much in an account paying 14% as he does in an account paying 5%. If he earns $152.75 in interest in one year from both accounts combined, how much did he invest altogether?
Total Principal in Both Accounts = $
an account pays 5% Assume the amount is $x ( 0.05)
Second account will be 3x @ 14% (0.14)
he earns $152.75 in interest in one year
Interest equation
0.05x + 0.14(3x) = 152.75
0.05x +0.42x =152.75
0.47x = 152.75
x = 152.78/0,47
x = 325.06
3x = 3*325.06 =975.18

Answer by ikleyn(52783)

.

Interest equation

    0.05x + 0.14(3x) = 152.75

    0.05x + 0.42x    = 152.75

       0.47x         = 152.75

           x         = 152.75/0.47

           x =   325
          3x = 3*325 = 975


Altogether  325 + 975 = 1300  dollars.    ANSWER

Solved.


Answer by greenestamps(13200)



Since he invested 3 times as much at 14% as at 5%, his weighted average percent return was

(3(14)+1(5))/4 = 11.75

Since he earned $152.75 in interest, the total amount invested was

$152.75/.1175 = $1300

ANSWER: $1300